An estuary has an average depth of 40.9 m. Estimate the length (km) of estuary required for it to experience tidal resonance (round to 0 decimal place) assuming diurnal tides.
Answer #1

Here the speed of the waves in the estuary is given by the formula :

{\displaystyle \scriptstyle {\sqrt {gh}}} where,

g is the acceleration of gravity and h is the depth of the estuary,

so, according to the question

depth = 40.9 m or 0.0409 km

g = 9.8 m/s2

usually for a continental shelf of depth 100 m the speed of long waves if approximately 30m/s

so if we calculate according to above conditions for a depth of 40.9 m the speed would be approximately 12.27m/s

So if the tidal period is 12 hours, a quarter wavelength shelf will have a width of about

= 40.9*12.27

= 501.843 m